OpenStudy (matt6288):

I need a math wizard! What is wrong with the dimensional analysis work in the following problem in which 5 feet/second was converted to miles per hour?

OpenStudy (matt6288):

5feet/second was converted to miles per hour? 5feet/1second*1mile/5280feet*1minute/60s… ≈ 0.0001 miles per hour.

OpenStudy (matt6288):

A. In the first conversation ratio, the second should be on top and the feet should be on bottom. B. In the second conversation ratio, the feet should be on top and the mile should be on bottom. C. In the third conversion ratio, the seconds should be on top and the minutes should be on the bottom. D. In the fourth conversation ratio, the hour should be on top and the minutes should be on bottom.

OpenStudy (amistre64):

\[\frac{ft}{sec}\frac{sec}{min}\frac{min}{hr}=\frac{ft}{hr}\] \[\frac{ft}{hr}\frac{mile}{ft}=\frac{mile}{hr}\]

OpenStudy (amistre64):

or some versions of that

OpenStudy (amistre64):

so the min/sec that they have is not a healthy operation

OpenStudy (matt6288):

so it would be A?

OpenStudy (amistre64):

hard to tell, that last bit needs to be an sec/hr i believe

OpenStudy (amistre64):

5feet 1mile 1minute sec feet sec is what they got so feet are off the top and we got mile/sec we need sec/hr to finish up

OpenStudy (amistre64):

the third conversion ... hard to count them, is whats bad at any rate

OpenStudy (amistre64):

is ft/sec the 1st conversion or the is it not counted as a conversion?

OpenStudy (amistre64):

ok, it reads as the first; so C is your best pick. oy that was torture :)

OpenStudy (matt6288):

Thx :)